OpenStudy (anonymous):

if 27^15-3^y, what is the value of y ?

OpenStudy (phi):

do you mean 27^15 = 3^y ? if so, you might know that 3*3*3= 27 or you could say 27= \( 3^3 \) replace 27 with \( 3^3 \) \[ \left(3^3\right)^{15} = 3^y\] now use the rule \[ \left(a^b\right)^c = a^{bc} \] to re-write the left side can you finish ?